1-1/(1+x)20-45%

solve for x
Answer #1

The value of x is computed as shown below:

1 - 1 / ( 1 + x ) 20 = 0.45

( 1 + x )20 - 1 = 0.45 ( 1 + x )20

0.55 ( 1 + x )20 = 1

(1 + x)20 = 1 / 0.55

(1 + x)20 = 1.818181818

1 + x = ( 1.818181818 ) 1 / 20

1 + x = 1.030343096

x = 0.030343096

Feel free to ask in case of any query relating to this question